# Improved Contact Graph Routing in Delay Tolerant Networks with Capacity and Buffer Constraints

### v1

@@ VIII Conclusions
In this work, we proposed enhanced contact-capacity management and buffer management to improve the state-of-the-art CGR routing algorithm. We proposed managing these resources by taking actions before or during the route search. Our solution is a global solution: It considers and tracks the evolution of the contact capacity and buffer occupation in all contacts and nodes of the network. The resulting route is optimal and respects contact capacity and node buffer constraints.
To manage contact capacity, we introduced contact splitting. It allows for consideration of the previously consumed portions of a contact by routed bundles. For the buffer management, we proposed temporary bundle-specific contact splitting and edge pruning to avoid any portion or succession of contacts that might result in buffer overflows. These temporary modifications depend on the forecast buffer tables that track the future states of the node buffer occupancy level as function of time.
We compared the performance of the proposed solution with the benchmark SR-CGR. The solution offers considerable gains in terms of time spent in the network and intermediate node complexity, measured as the number of rerouting calls.
Limitations. Although the proposed solution offers improvements over the state-of-the-art while meeting capacity and buffer constraints, it has some limitations that could be taken into account for further improvements. Our solution is based on SR-CGR, which is usually used to reduce CGR complexity. This reduction is achieved by performing one route search per bundle rather than multiple route searches (per hop routing). Nevertheless, our solution splits a contact into up to two contacts, when a bundle occupies a contact capacity or a buffer space. The splitting action increases the size of the CP, which directly impacts the complexity of the route search. In our solution, the additional complexity is tractable because as the route search is performed only on source nodes and not on all low-power intermediate nodes. Nevertheless, when moving to a distributed solution, one should consider such additional complexity. The centralized solution is also a limitation of this study. The examination of an efficient distributed solution is therefore an interesting research direction as improvement of this work.
Future work. We plan to assess proposed cislunar constellations, from a communication perspective, using the presented routing technology. We believe that the proposed routing algorithm may help the design of the communication and storage systems to be embedded in the satellites of such lunar constellations.
We also plan to investigate traffic-level solutions rather than a per-bundle solution. In the state-of-the-art as in this work, the route is chosen based on optimal criterion at the bundle level, not for the entire traffic to be transmitted. Therefore, this per-bundle strategy is likely to be not optimal for the overall communication performance.
Finally, as mentioned in the previous “limitations” subsection, investigating distributed solutions rather than a centralized one is also of interest.

### v2

@@ VIII Conclusions
In this work, we proposed enhanced contact capacity and buffer management to improve the state-of-the-art CGR routing algorithm. We proposed managing these resources by taking actions before or during the route search. Our solution is a global solution: It considers and tracks the evolution of the contact capacity and buffer occupation in all contacts and nodes of the network. The resulting route is optimal and respects contact capacity and node buffer constraints.
To manage contact capacity, we introduced contact splitting. It enables tracking of contact portions previously consumed by routed bundles. For the buffer management, we proposed two techniques: temporary bundle-specific contact splitting and edge pruning. These techniques help avoid any portion or succession of contacts that might result in buffer overflows. These temporary modifications depend on the forecast buffer tables that track the future states of the node buffer occupancy levels as a function of time.
We compared the performance of the proposed solution with the benchmark SR-CGR. Our solution offers considerable gains in terms of time spent in the network and intermediate node complexity, measured as the number of rerouting calls. It is worth mentioning that the increase in CP size due to our splitting actions, and thus the route search complexity, is relatively small.
Limitations. Although the proposed solution offers improvements over the state-of-the-art while meeting capacity and buffer constraints, our solution remains centralized. The examination of an efficient distributed solution is therefore an interesting research direction for improvement of this work.
Future work. We plan to assess proposed cislunar constellations, from a communication perspective, using the presented routing technology. We believe that the proposed routing algorithm may help with the design of the communication and storage systems to be embedded in the satellites of such lunar constellations.
We also plan to investigate traffic-level solutions rather than a per-bundle solution. In the state-of-the-art, like in this work, the route is chosen based on optimal criteria at the bundle level, not for the entire traffic to be transmitted. Therefore, this per-bundle strategy is unlikely to be optimal for the overall communication performance. Finally, as mentioned in the previous “limitations” subsection, investigating distributed solutions rather than a centralized one is also of interest.